OpenStudy (anonymous):

The non-profit organization you volunteer for is throwing a fundraiser cook out. You are in charge of buying the hamburgers, which cost $3 per pound and hotdogs, which cost $2 per pound. The meat budget you are given totals $600. The inequality 3x + 2y ≤ 600 represents the possible combinations of pounds of hamburgers (x) and hotdogs (y) you can buy. a. 200 pounds of hamburgers and 140 pounds of hotdogs b. 150 pounds of hamburgers and 60 pounds of hotdogs c. 100 pounds of hamburgers and 240 pounds of hotdogs d. 240 pounds of hamburgers and 40 pounds of hotdogs

OpenStudy (jdoe0001):

well, this case is the same as the one before, you have to do testing of the values in the inequality you're given 2 points really, hamburgers = x hotdogs = y like say let's test the first one A) you're given the x, y set of (200, 140) let's check that \(\bf 3x + 2y \le 600\\ 3(200)+2(140) \le 600 \implies 600+280 \le 600 \implies 880 \le 600\) well we know that \(\bf 880 \cancel{\le} 600\) so that's not it check B

OpenStudy (texaschic101):

she's not online

OpenStudy (jdoe0001):

yes, I notice that, thus I left just the 1st test :)

OpenStudy (texaschic101):

good idea